# Superquantile-Gibbs Relaxation for Minima-selection in Bilevel Optimization

### v2

@@ 1.4 Related work
Beyond such regularity-based assumptions (strong convexity or global PŁ), recent works proposed alternative formulations of BLO that target nonconvex lower-level problems. For instance, bolte2025bilevel worked under a (piecewise) Morse-type parametric qualification, showed that the graphs of lower-level critical points and local minima generically decomposed into finitely many smooth manifolds, and used this structure to define branch-wise hyper-objectives and study gradient methods that updated the upper-level variable along individual branches. arbel2022non instead introduced bilevel games with selection maps, where a selection map chose a specific critical point (for instance, the limit of a gradient flow) for each upper-level parameter, and showed that standard bilevel algorithms based on truncated unrolling and implicit differentiation tended to approximate such selection-based games rather than the classical bilevel formulation. jiang2025correspondence further argued that, for general nonconvex lower levels, the classical hyper-objective may be ill-posed when global minimizers are unattainable; they proposed a correspondence-driven hyper-objective that evaluated the upper-level loss at the output of a prescribed lower-level algorithm, applied Gaussian smoothing to this mapping, and developed a projected stochastic-gradient method with convergence and oracle-complexity guarantees for the smoothed problem. In contrast, our work retains the classical (pessimistic or optimistic) hyper-objective, combines it with a local PŁ condition to obtain a manifold structure for the lower-level minimizers, and uses a superquantile–Gibbs relaxation to define an explicit selection rule together with dimension-aware complexity bounds.
@@
The approaches discussed above treat lower-level non-uniqueness either by enforcing regularity conditions (strong convexity or global PŁ) or by embedding explicit selection mechanisms (Morse branches, selection maps, or algorithmic correspondences), and thus remain essentially in an optimistic or selection-based setting. By contrast, the pessimistic formulation evaluates each upper-level decision against the worst upper-level value over all lower-level minimizers, which is natural when the follower may act adversarially or unreliably (beck2023survey). This risk-averse viewpoint was studied in mathematical programming, including structural and complexity results for general pessimistic models (wiesemann2013pessimistic), surveys of modelling and solution schemes (liu2018survey), and necessary optimality conditions based on generalized differentiation and value-function techniques (dempe2014necessary). On the algorithmic side, several works constructed single-level relaxations of the pessimistic problem: zeng2020practical designed a tight relaxation and a computation scheme for (mainly linear) pessimistic bilevel problems; benchoouk2025scholtes analyzed a Scholtes-type relaxation of the KKT system; and antoniou2024delta proposed a $\delta$-perturbed formulation that enforced an approximate unique follower response and derived error bounds between the perturbed and original pessimistic problems. Pessimistic ideas also appeared in machine learning: ustun2024pessimistic formulated hyperparameter tuning as a pessimistic bilevel problem over a finite set of tasks, and jiang2021learningdefend interpreted adversarial training as a bilevel game in which a learned attacker played the role of the follower. All these approaches, however, either (i) resolved non-uniqueness by modifying the lower-level problem (e.g., $\delta$-perturbations, KKT relaxations) or (ii) specialized to finite or discrete follower responses, and gradient-based methods with nonasymptotic guarantees rarely addressed the continuous pessimistic setting with a non-singleton solution set. In contrast, we work directly with the pessimistic hyper-objective $F_{\max}$ under a local PŁ condition, model pessimism via a superquantile risk over Gibbs samples on the minimizer manifolds, and obtain dimension-aware complexity bounds that explicitly depend on the intrinsic dimension k of $\mathcal{S}(\theta)$ rather than on the ambient dimension.
@@ 2 Summary of Assumptions and Preliminaries
In this work, we focus on the setting where the lower-level objective is coercive and satisfies local Polyak-Łojasiewicz () condition for all upper-level variable $\theta$, formally stated as follows.
@@
    $g$ is a $\mathcal{C}^{2}$ and locally PŁ function, and $g$ is a $\mathcal{C}^{4}$ function in every neighborhood $\mathcal{N}(\mathcal{M}^{\prime})$555gong2024poincare require $g\in\mathcal{C}^{3}$. Here we slightly enhance the assumption to $\mathcal{C}^{4}$ to enable Section 2.1.2..
@@
Proposition 2.3 (Connectivity, Proposition 3 in (gong2024poincare)).
@@
Consequently, all local minima of a function are global minima which are all connected. Further, rebjock2023fast proved the local manifold structure for $\mathcal{M}$. It is later extend to a global version in (gong2024poincare), where they also provide an explicit statement about the boundary.
@@
The second condition in Section˜2 is used solely to establish the connectedness of the local minima set via the mountain pass theorem (see the proof of Proposition 3 in (gong2024poincare) for details). Our results remain valid if this condition is replaced with the assumption that $\mathcal{S}(\theta)$ is connected. Empirically, such connectedness has been widely observed in over-parameterized deep learning settings; see, e.g., (draxler2018essentially; nguyen2019connected; kuditipudi2019explaining).
Proposition 2.5 (Manifold structure, Corollary 1 in (gong2024poincare)).
@@
Following (gong2024poincare), we make the next assumption to exclude the manifolds with ill-conditioned chart. We also use it to establish a nonzero lower bound on the injectivity radius, allowing us to analyze the volume-radius relationship of geodesic balls, which is crucial for our approximation results.
@@
The following statements hold for all $\theta\in\Theta$. $\bullet$ $f$ is a $\mathcal{C}^{1}$ function and $L_{f,1}$-Lipschitz, and $L_{f,2}$-smooth w.r.t. $(\theta,x)$. For all $x\in\mathbb{R}^{d}$, $|f(\theta,x)|\leq C_{f}\|x\|^{\mathsf{n}_{f}}+D_{f}$ where $C_{f},D_{f}\geq 0$ are two constant and $\mathsf{n}_{f}\geq 1$ is an integer constant. $\bullet$ $g$ is a $\mathcal{C}^{2}$ function and $L_{g,2}$-smooth w.r.t. $\theta$ for every $x\in\mathbb{R}^{d}$. For all $x\in\mathbb{R}^{d}$, $\|\partial_{\theta}g(\theta,x)\|\leq C_{g}\|x\|^{\mathsf{n}_{g}}+D_{g}$ where $C_{g},D_{g}\geq 0$ are two constant and $\mathsf{n}_{g}\geq 1$ is an integer constant. $\bullet$ The Hessian $\partial_{x}^{2}g(\theta,x)$ is continuous w.r.t. $(\theta,x)$. $\bullet$ Beyond a compact set, $g(\theta,\cdot)$ satisfies the quadratic growth, i.e., $\exists\mathsf{D}>0$ such that $\forall|x|\geq\mathsf{D}$, $\mu_{\mathsf{qg}}\cdot\mathrm{dist}^{2}(x,\mathcal{S}(\theta))/2\leq g(\theta,x)-\min_{z\in\mathbb{R}^{d}}g(\theta,z)$ [^6] W.L.O.G., we assume that $\mathsf{D}$ is sufficiently large so that $\mathcal{S}(\theta)\subseteq\mathbb{B}_{d}(0;\mathsf{D})$ for all $\theta\in\Theta$. This assumption implies the coercivity of function $g(\theta,\cdot)$..
The regularity conditions in Section˜2, such as Lipschitzness and smoothness, are standard in the literature (kwon2023penalty; chen2023bilevel). However, unlike kwon2023penalty, we do not assume a bounded $x$-domain in the lower-level problem. Instead, we impose quadratic growth beyond a compact set, ensuring $\mu_{g}^{\lambda}$ is well-defined (hasenpflug2024wasserstein). Additionally, while kwon2023penalty assumes boundedness of $f$, we relax this to a polynomial bound condition. We also assume a bounded $\theta$-domain and polynomial bound on $\partial_{\theta}g$, which we plan to further relax in future work.
@@ 2.1.1 Implications of the Embedding Submanifold Structure in Section˜2
Following Remark 1 in (gong2024poincare), WLOG and for the ease of notation, we assume that the atlas of the manifold $\mathcal{S}$ consists of a single chart $(\mathcal{U},\phi)$, where $\mathcal{U}=\mathcal{S}$. Let $\Gamma=\phi(\mathcal{U})\subseteq\mathbb{R}^{\texttt{k}}$ denote the coordinate domain [^7] The standard technique of partition of unity can be used to translate the results to the multi-chart case.. The corresponding parametrization (the inverse of the chart) can be represented as $x=\mathsf{S}(u):=[m_{1}(u),\ldots,m_{d}(u)]^{\top}$ for all $u\in\Gamma$, where $m_{j}:\Gamma\to\mathbb{R}$ for $j\in[d]$ are $\mathcal{C}^{2}$ coordinate functions.
@@
For $x\in\mathcal{S}$, the tangent space at $x$ is denoted by $T_{x}(\mathcal{S})$. The orthogonal complement of $T_{x}(\mathcal{S})$ in $\mathbb{R}^{d}$ is denoted by $T_{x}(\mathcal{S})^{\perp}$. The tubular neighborhood $\mathcal{N}(\varepsilon)$ of $\mathcal{S}$ for $\varepsilon>0$ is defined as: $\mathcal{N}(\varepsilon):=\{x+v:x\in\mathcal{S},v\in T_{x}(\mathcal{S})^{\perp},\|v\|<\varepsilon\}.$ A key property of the $\mathcal{C}^{2}$ embedded submanifold is as follows [^8] For simplicity, we assume a single chart and omit the dependence on the inclusion $i_{\mathcal{S}}$. (lee2024introduction, Theorem 6.17): There exists $\varepsilon_{0}>0$ such that for all $0<\varepsilon\leq\varepsilon_{0}$, there is a $\mathcal{C}^{2}$ diffeomorphism $\Phi$:
@@
where $\mathbb{H}^{1}(\mu)$ denotes the Sobolev space weighted by $\mu$. Importantly, for all functions $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$, the constant $C_{PI}$ for the Gibbs probability measure $\mu^{\lambda}\propto\exp(-g/\lambda)$ is independent of $\lambda$ for all sufficiently small $\lambda\leq\lambda_{0}$, where both $\lambda_{0}$ and an upper bound of $C_{PI}$ are explicitly given in (gong2024poincare).
@@ 2.1.2 Limiting measure of Gibbs measures
To sample approximately from the optimal set $\mathcal{S}$, we select a temperature parameter $\lambda$ approaching zero, which causes the Gibbs measure (3) to concentrate on $\mathcal{S}$. Let $\mu^{0}:=\lim_{\lambda\to 0}\mu^{\lambda}$ denote the weak limit of the Gibbs measures. By combining the manifold characterization of the optimal set for functions in (gong2024poincare) with the asymptotic analysis in (hwang1980laplace, Theorem 3.1), the following lemma establishes that $\mu^{0}$ is well-defined, has support restricted to $\mathcal{S}$, and admits a closed-form density function with respect to the Riemannian volume measure $d\mathcal{M}$ on the manifold.
Proposition 2.10 (Theorem 3.1 in (hwang1980laplace)).
@@
The following proposition from (hasenpflug2024wasserstein, Theorem 3.6) proves the non-asymptotic convergence of $\mu^{\lambda}$ as $\lambda\to 0$ to $\mu^{0}$ (defined above) in the Wasserstein-1 sense [^11] We use the assumption on $g$ in Section 2 to ensure that the conditions in (hasenpflug2024wasserstein, Theorem 3.6) hold. Further discussion on the validity of their conditions can be found in Appendix A.2.. Establishing convergence in $\mathbb{W}_{1}$ is crucial for our analysis because it directly bounds the approximation error between the superquantile-Gibbs surrogate and the superquantile loss taken over the limiting measure $\mu^{0}$ (to be defined in (11)). This bound is established by exploiting the Lipschitz continuity of the upper-level objective (see Section˜4.2).
Proposition 2.11 (Theorem 3.6 in (hasenpflug2024wasserstein) ).
@@ 3.1 Lipschitz Continuity of the Solution Mapping
Since the condition provides only a local PŁ inequality in a neighborhood of the minimizer set $\mathcal{S}(\theta)$, Lipschitz stability of the solution mapping does not follow from standard global error-bound arguments. By contrast, under a global PŁ condition for $g(\theta,\cdot)$, one obtains a global error bound and then deduces Lipschitz continuity of $\theta\mapsto\mathcal{S}(\theta)$ directly from the smoothness of $g$ on $\theta$ (see, e.g., (kwon2023penalty; chen2023bilevel)). In our local setting, the obstruction is geometric: a local error bound alone still allows the minimizer set to have multiple disconnected components, and minimizers could jump from one component to another under arbitrarily small changes in $\theta$, precluding any Hausdorff Lipschitz control. This is why we need that $\mathcal{S}(\theta)$ is connected under our assumptions (see Section˜2) to prove a quantitative stability result showing that minimizers cannot jump and must move continuously under small perturbations of $\theta$. The next lemma makes this precise.
@@
By the local PŁ condition and $\mathcal{C}^{2}$ regularity, $g(\theta,\cdot)$ satisfies a local error bound in a smaller neighborhood of $\mathcal{S}(\theta)$ (rebjock2023fast, Proposition 2.2): there exists an open set $\mathcal{U}(\theta)\subseteq\mathcal{N}(\mathcal{S}(\theta))$ with $\mathcal{S}(\theta)\subset\mathcal{U}(\theta)$ such that
@@
Having established the local Lipschitz continuity of $\mathcal{S}(\theta)$ in Lemma 3.1, we now turn to the regularity of the hyper-objective $F_{\max}$. The following corollary follows by combining the local Lipschitz continuity of $\theta\mapsto\mathcal{S}(\theta)$ with the Lipschitz continuity of $f(\theta,x)$ [^12] A standard chaining argument shows that if $|F_{\max}(\theta_{1})-F_{\max}(\theta_{2})|\leq L\|\theta_{1}-\theta_{2}\|$ whenever $\|\theta_{1}-\theta_{2}\|\leq\gamma$, then the same constant works uniformly on $\Theta$.; see, e.g., (chen2023bilevel, Theorem 3.1).

### v4

@@ 1.4 Related Work
Beyond such regularity-based assumptions (strong convexity or global PŁ), recent works proposed alternative formulations of BLO that target nonconvex lower-level problems. For instance, Bolte et al. [11] worked under a (piecewise) Morse-type parametric qualification, showed that the graphs of lower-level critical points and local minima are generically decomposed into finitely many smooth manifolds, and used this structure to define branch-wise hyper-objectives and study gradient methods that updated the upper-level variable along individual branches. Arbel and Mairal [4] introduced bilevel games with selection maps, where the selection map chooses a specific lower-level critical point for each upper-level parameter (e.g., the limit of a gradient flow from a given initialization). They showed that popular unrolled/iterative-differentiation bilevel algorithms with a finite inner budget approximate equilibria of these selection-based games up to approximation errors, and they extended implicit differentiation to characterize the selection and proposed a correction that removes the finite-budget error. Jiang et al. [43] further argued that, for general nonconvex lower levels, the classical hyper-objective may be ill-posed when global minimizers are unattainable; they proposed a correspondence-driven hyper-objective that evaluated the upper-level loss at the output of a prescribed lower-level algorithm, applied Gaussian smoothing to this mapping, and developed a projected stochastic gradient method with oracle-complexity guarantees for the smoothed problem. Another recent line develops Moreau-envelope single-level reformulations as an alternative to classical value-function and KKT-based complementarity-constrained reductions. Bai et al. [5] derive directional necessary optimality conditions in this framework, Gao et al. [30] propose a difference-of-weakly-convex reformulation with a sequentially convergent algorithm, and Lu [65] builds on these ideas to obtain a two-sided smoothed primal–dual method for general nonconvex bilevel problems with KKT-type convergence guarantees. In contrast, our work retains the classical (pessimistic or optimistic) hyper-objective, combines it with a local PŁ condition to obtain a manifold structure for the lower-level minimizers, and uses a superquantile–Gibbs relaxation to convert the minima-selection into sampling from a Gibbs measure.
@@
The approaches discussed above treat lower-level non-uniqueness either by enforcing regularity conditions (strong convexity or global PŁ) or by embedding explicit selection mechanisms (Morse branches, selection maps, or algorithmic correspondences), and thus remain essentially in an optimistic or selection-based setting. By contrast, the pessimistic formulation evaluates each upper-level decision against the worst upper-level value over all lower-level minimizers, which is natural when the follower may act adversarially or unreliably [6]. This risk-averse viewpoint was studied in mathematical programming, including structural and complexity results for general pessimistic models [82], surveys of modelling and solution schemes [60], and necessary optimality conditions based on generalized differentiation and value-function techniques [24]. On the algorithmic side, several works constructed single-level relaxations of the pessimistic problem: Zeng and Ye [92] designed a tight relaxation and a computation scheme for (mainly linear) pessimistic bilevel problems; Benchoouk et al. [7] analyzed a Scholtes-type relaxation of the KKT system; and Antoniou et al. [2] proposed a $\delta$-perturbed formulation that enforced an approximate unique follower response and derived error bounds between the perturbed and original pessimistic problems. Pessimistic ideas also appeared in machine learning: Ustun et al. [79] formulated hyperparameter tuning as a pessimistic bilevel problem over a finite set of tasks and highlighted advantages of this formulation over the optimistic counterpart, and Jiang et al. [42] interpreted adversarial training as a bilevel game in which a learned attacker played the role of the follower. All these approaches, however, either (i) resolved non-uniqueness by modifying the lower-level problem (e.g., $\delta$-perturbations, KKT relaxations) or (ii) specialized to finite or discrete follower responses, rarely addressed the continuous pessimistic setting with a non-singleton solution set. In contrast, we work directly with the pessimistic hyper-objective $F_{\max}$ under a local PŁ condition, model pessimism via a superquantile risk over Gibbs samples on the minimizer manifolds, and obtain complexity results that explicitly depend on the intrinsic dimension k of $\mathcal{S}(\theta)$ rather than on the ambient dimension.
@@ 2 Summary of Assumptions and Preliminaries
In this work, we focus on the setting where the lower-level objective is coercive and satisfies a local Polyak-Łojasiewicz () condition for all upper-level variables $\theta\in\Theta$, formally stated as follows.
@@
    $g$ is a $\mathcal{C}^{2}$ and locally PŁ function, and $g$ is a $\mathcal{C}^{4}$ function in every neighborhood $\mathcal{N}(\mathcal{M}^{\prime})$ [^5] Gong et al. [35] require $g\in\mathcal{C}^{3}$. Here we slightly enhance the assumption to $\mathcal{C}^{4}$ to enable Section 2.1.2..
@@
Proposition 2.3 (Connectivity, Proposition 3 in [35]).
@@
Consequently, all local minima of a function are global minima which are all connected. Further, Rebjock and Boumal [73] proved the local manifold structure of $\mathcal{M}$. This was later extend to a global version in [35], where they also provide an explicit statement about the boundary.
@@
The second condition in Section˜2 is used solely to establish the connectedness of the local minima set via the mountain pass theorem (see the proof of Proposition 3 in [35] for details). Our results remain valid if this condition is replaced with the assumption that $\mathcal{S}(\theta)$ is connected. Empirically, such connectedness has been widely observed in over-parameterized deep learning settings; see, e.g., [27, 69, 48].
Proposition 2.5 (Manifold structure, Corollary 1 in [35]).
@@
Following [35], we make the next assumption to exclude the manifolds with ill-conditioned charts. We also use it to establish a nonzero lower bound on the injectivity radius, allowing us to analyze the volume-radius relationship of geodesic balls, which is crucial for our approximation results.
@@
The following statements hold for all $\theta\in\Theta$. $\bullet$ $f$ is a $\mathcal{C}^{1}$ function and $L_{f,1}$-Lipschitz, and $L_{f,2}$-smooth w.r.t. $(\theta,x)$. For all $x\in\mathbb{R}^{d}$, $|f(\theta,x)|\leq C_{f}\|x\|^{\mathsf{n}_{f}}+D_{f}$ where $C_{f},D_{f}\geq 0$ are two constants and $\mathsf{n}_{f}\geq 1$ is an integer constant. $\bullet$ $g$ is a $\mathcal{C}^{2}$ function and $L_{g,2}$-smooth w.r.t. $\theta$ for every $x\in\mathbb{R}^{d}$. For all $x\in\mathbb{R}^{d}$, $\|\partial_{\theta}g(\theta,x)\|\leq C_{g}\|x\|^{\mathsf{n}_{g}}+D_{g}$ where $C_{g},D_{g}\geq 0$ are two constants and $\mathsf{n}_{g}\geq 1$ is an integer constant. $\bullet$ The Hessian $\partial_{x}^{2}g(\theta,x)$ is continuous w.r.t. $(\theta,x)$. $\bullet$ Beyond a compact set, $g(\theta,\cdot)$ satisfies the quadratic growth, i.e., $\exists\mathsf{D}>0$ such that $\forall|x|\geq\mathsf{D}$, $\mu_{\mathsf{qg}}\cdot\mathrm{dist}^{2}(x,\mathcal{S}(\theta))/2\leq g(\theta,x)-\min_{z\in\mathbb{R}^{d}}g(\theta,z)$ [^6] W.L.O.G., we assume that $\mathsf{D}$ is sufficiently large so that $\mathcal{S}(\theta)\subseteq\mathbb{B}_{d}(0;\mathsf{D})$ for all $\theta\in\Theta$. This assumption implies the coercivity of function $g(\theta,\cdot)$..
The regularity conditions in Section˜2, such as Lipschitzness and smoothness, are standard in the literature [50, 14]. However, unlike in [50], we do not assume a bounded $x$-domain in the lower-level problem. Instead, we impose quadratic growth beyond a compact set, ensuring $\mu_{g}^{\lambda}$ is well-defined [38]. Additionally, whereas [50] assumes boundedness of $f$, we relax this to a polynomial bound condition. We also assume a bounded $\theta$-domain and polynomial bound on $\partial_{\theta}g$, which we plan to further relax in future work.
@@ 2.1.1 Implications of the Embedding Submanifold Structure in Section˜2
Following Remark 1 in [35], WLOG and for the ease of notation, we assume that the atlas of the manifold $\mathcal{S}$ consists of a single chart $(\mathcal{U},\phi)$, where $\mathcal{U}=\mathcal{S}$. Let $\Gamma=\phi(\mathcal{U})\subseteq\mathbb{R}^{\texttt{k}}$ denote the coordinate domain [^7] The standard technique of partition of unity can be used to translate the results to the multi-chart case.. The corresponding parametrization (the inverse of the chart) can be represented as $x=\mathsf{S}(u):=[m_{1}(u),\ldots,m_{d}(u)]^{\top}$ for all $u\in\Gamma$, where $m_{j}:\Gamma\to\mathbb{R}$ for $j\in[d]$ are $\mathcal{C}^{2}$ coordinate functions.
@@
For $x\in\mathcal{S}$, the tangent space at $x$ is denoted by $T_{x}(\mathcal{S})$. The orthogonal complement of $T_{x}(\mathcal{S})$ in $\mathbb{R}^{d}$ is denoted by $T_{x}(\mathcal{S})^{\perp}$. The tubular neighborhood $\mathcal{N}(\varepsilon)$ of $\mathcal{S}$ for $\varepsilon>0$ is defined as: $\mathcal{N}(\varepsilon):=\{x+v:x\in\mathcal{S},v\in T_{x}(\mathcal{S})^{\perp},\|v\|<\varepsilon\}.$ A key property of the $\mathcal{C}^{2}$ embedded submanifold is as follows [^8] For simplicity, we assume a single chart and omit the dependence on the inclusion $i_{\mathcal{S}}$. [52, Theorem 6.24]: There exists $\varepsilon_{0}>0$ such that for all $0<\varepsilon\leq\varepsilon_{0}$, there is a $\mathcal{C}^{2}$ diffeomorphism $\Phi$:
@@
where $\mathbb{H}^{1}(\mu)$ denotes the Sobolev space weighted by $\mu$. Importantly, for all functions $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$, the constant $C_{PI}$ for the Gibbs probability measure $\mu^{\lambda}\propto\exp(-g/\lambda)$ is independent of $\lambda$ for all sufficiently small $\lambda\leq\lambda_{0}$, where both $\lambda_{0}$ and an upper bound of $C_{PI}$ are explicitly given in [35].
@@ 2.1.2 Limiting Measure of Gibbs Measures
To sample approximately from the optimal set $\mathcal{S}$, we select a temperature parameter $\lambda$ approaching zero, which causes the Gibbs measure (3) to concentrate on $\mathcal{S}$. Let $\mu^{0}:=\lim_{\lambda\to 0}\mu^{\lambda}$ denote the weak limit of the Gibbs measures. By combining the manifold characterization of the optimal set for functions in [35] with the asymptotic analysis in [41, Theorem 3.1], the following lemma establishes that $\mu^{0}$ is well-defined, has support restricted to $\mathcal{S}$, and admits a closed-form density function with respect to the Riemannian volume measure $d\mathcal{M}$ on the manifold.
Proposition 2.10 (Theorem 3.1 in [41]).
@@
The following proposition from [38, Theorem 3.6] proves the non-asymptotic convergence of $\mu^{\lambda}$ as $\lambda\to 0$ to $\mu^{0}$ (defined above) in the Wasserstein-1 sense [^11] We use the assumption on $g$ in Section 2 to ensure that the conditions in [38, Theorem 3.6] hold. Further discussion on the validity of their conditions can be found in Appendix A.3.. Establishing convergence in $\mathbb{W}_{1}$ is crucial for our analysis because it directly bounds the approximation error between the superquantile-Gibbs surrogate and the superquantile loss taken over the limiting measure $\mu^{0}$ (to be defined in (11)). This bound is established by exploiting the Lipschitz continuity of the upper-level objective (see Section˜4.2).
Proposition 2.11 (Theorem 3.6 in [38] ).
@@ 3.1 Lipschitz Continuity of the Solution Mapping
Since the condition provides only a local PŁ inequality in a neighborhood of the minimizer set $\mathcal{S}(\theta)$, Lipschitz stability of the solution mapping does not follow from standard global error-bound arguments. By contrast, under a global PŁ condition for $g(\theta,\cdot)$, one obtains a global error bound and then deduces Lipschitz continuity of $\theta\mapsto\mathcal{S}(\theta)$ directly from the smoothness of $g$ on $\theta$ (see, e.g., [50, 14]). In our local setting, the obstruction is geometric: a local error bound alone still allows the minimizer set to have multiple disconnected components, and minimizers could jump from one component to another under arbitrarily small changes in $\theta$, precluding any Hausdorff Lipschitz control. This is why we need that $\mathcal{S}(\theta)$ is connected under our assumptions (see Section˜2) to prove a quantitative stability result showing that minimizers cannot jump and must move continuously under small perturbations of $\theta$. The next lemma makes this precise.
@@
By the local PŁ condition and $\mathcal{C}^{2}$ regularity, $g(\theta,\cdot)$ satisfies a local error bound in a smaller neighborhood of $\mathcal{S}(\theta)$ [73, Propositions 2.2 & 2.8]: there exists an open set $\mathcal{U}(\theta)\subseteq\mathcal{N}(\mathcal{S}(\theta))$ with $\mathcal{S}(\theta)\subset\mathcal{U}(\theta)$ such that
@@
Having established the local Lipschitz continuity of $\mathcal{S}(\theta)$ in Lemma 3.1, we now turn to the regularity of the hyper-objective $F_{\max}$. The following corollary follows by combining the local Lipschitz continuity of $\theta\mapsto\mathcal{S}(\theta)$ with the Lipschitz continuity of $f(\theta,x)$ [^12] A standard chaining argument shows that if $|F_{\max}(\theta_{1})-F_{\max}(\theta_{2})|\leq L\|\theta_{1}-\theta_{2}\|$ whenever $\|\theta_{1}-\theta_{2}\|\leq\gamma$, then the same constant works uniformly on $\Theta$.; see, e.g., [14, Theorem 3.1].